Please solve the provided system of linear equations by the substitution method.
6x+9y=6
2x=2-3y

Answer by TimothyLamb(4379)

6x + 9y = 6
x + (3/2)y = 1
x = 1 - (3/2)y
---
2x = 2 - 3y
---
2x = 2 -3y
2(1 - (3/2)y) = 2 - 3y
2 - 3y = 2 - 3y
2 = 2
---
the two linear equations appear to describe the same line
---
check:
put in slope-intercept form:
9y = -6x + 6
y = (-2/3)x + 2/3
---
put in slope-intercept form:
2x = 2 - 3y
3y = -2x + 2
y = (-2/3)x + 2/3
---
Answer:
the two linear equations describe the same line
---
Solve and graph linear equations:
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php